OpenStudy (anonymous):

Can anyone explain to me how to differentiate 10^x .. the derivative is 10^x*ln(10)

OpenStudy (anonymous):

change the 10 to e^ln10 and then use the chain rule so the derivative of e^(ln10*x) is e^(ln10*x) * ln10 =10^x * ln (10)

OpenStudy (anonymous):

Awesome, thanks.

myininaya (myininaya):

\[y=10^x \] Take natural log of both sides \[\ln(y)=\ln(10^x)\] Using properties of log namely ln(x^r)=rln(x)<-this law we can rewrite as \[\ln(y)=x \ln(10)\] Now differentiating both sides we receive \[\frac{y'}{y}= \ln(10) \text{ note: remember (xc)'=(cx)'=c(x)'=c(1)=c (c is constant))}\] multipliing both sides by y we receive \[y'=y \ln(10)\] but like omg y was 10^x so \[y'=10^x \ln(10)\]

OpenStudy (amistre64):

id say to memorize a general rule for any base. [B^x]' = B^x ln(B) when the Base is "e" we get: [e^x]' = e^x ln(e) , and since ln(e) = 1 it can be written as: = e^x

OpenStudy (anonymous):

hey myininaya, how did you get words to show up not stupid in your equation?

OpenStudy (amistre64):

\text{words}

OpenStudy (anonymous):

hmm, learn something new everyday.